Assume that a sample is used to estimate a population proportion μμ. Find the margin of error M.E. that corresponds to a sample of size 627 with a mean of 24.4 and a standard deviation of 5.7 at a confidence level of 98%.

Report ME accurate to one decimal place because the sample statistics are presented with this accuracy.
M.E. =

Answer should be obtained without any preliminary rounding. However, the critical value may be rounded to 3 decimal places.
Answer #1

Answer:

Given,

Mean = xbar = 24.4

standard deviation = s = 5.7

n = 627

alpha = 1 - 0.98

= 0.02

degree of freedom = n - 1

= 627 - 1

= 626

Critical value = t (alpha/2 , df) = t (0.01 , 626) = 2.332

Margin of error = E = t*s/sqrt(n)

substitute values

= 2.332*5.7 / sqrt(627)

E = 0.5

Now 98% confidence interval ,

= xbar +/- E

substitute values

= 24.4 +/- 0.5

= (24.4 - 0.5 , 24.4 + 0.5)

= (23.9 , 24.9)